A ladder is placed against the outside of a house at the roof line. The ladder is 10 feet long. The bottom of the ladder is plaed 3 feet from the house. What is the height of the side of the house? Round off answer to the nearest tenth if necessary
Answer by nyc_function(2741)

Let h = approximate height of building
h^2 + 3^2 = 10^2
h^2 + 9 = 100
h^2 = 100 - 9
h^2 = 91
sqrt(h^2) = sqrt{91}
h is approximately 9.5
nycfunction@yahoo.com
Guido